# Zero-Shot Cross-Lingual Document-Level Event Causality Identification with Heterogeneous Graph Contrastive Transfer Learning

### 05 Mar 2024 v1

@@ Abstract
Event Causality Identification (ECI) refers to detect causal relations between events in texts. However, most existing studies focus on sentence-level ECI with high-resource language, leaving more challenging document-level ECI (DECI) with low-resource languages under-explored. In this paper, we propose a Heterogeneous Graph Interaction Model with Multi-granularity Contrastive Transfer Learning (GIMC) for zero-shot cross-lingual document-level ECI. Specifically, we introduce a heterogeneous graph interaction network to model the long-distance dependencies between events that are scattered over document. Then, to improve cross-lingual transferability of causal knowledge learned from source language, we propose a multi-granularity contrastive transfer learning module to align the causal representations across languages. Extensive experiments show our framework outperforms previous state-of-the-art model by 9.4% and 8.2% of average F1 score on monolingual and multilingual scenarios respectively. Notably, in multilingual scenario, our zero-shot framework even exceeds GPT-3.5 with few-shot learning by 24.3% in overall performance.
Keywords: zero-shot cross-lingual, event causality identification, graph contrastive learning
@@
${}^{3}$ Ant Group, Hangzhou, China
@@ 2.2.   Heterogeneous Graph Interaction Network
After construct heterogeneous graph, we apply Graph Attention Netstudies v2 (GATv2) Brody et al. (2022) to model the global interactions. For each edge ($j,i$), the scoring function $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ indicates the importance of the features of the neighbor $j\in\mathcal{N}_{i}$ to the node $i$:
@@ 2.4.   Training
In our implementations, we employ the base versions of the language-specific PLMs, i.e., BotXO2 for Danish, BERT for English, BETO for Spanish, BERTurk for Turkish, UrduHack for Urdu and the multilingual PLMs, i.e., mBERT and XLMR as backbone respectively. The learning rate is initialized as 1e-3 with a linear decay. We use the AdamW algorithm Loshchilov and Hutter (2017) to optimize model parameters. The batch size is set to 1, the number of GATv2 layers is 3. The number of training epochs is 60. Each experiment is conducted on NVIDIA GeForce RTX 3090 GPUs.
@@ 3.3.   Baselines
(3) RichGCNTran Phu and Nguyen (2021) implement several interaction graphs, by learning a linear combination of the adjacency matrices of these graphs to get a final graph and uses GCN to capture capture relevant context information.

### 22 Mar 2024 v2

@@ Abstract
Event Causality Identification (ECI) refers to the detection of causal relations between events in texts. However, most existing studies focus on sentence-level ECI with high-resource languages, leaving more challenging document-level ECI (DECI) with low-resource languages under-explored. In this paper, we propose a Heterogeneous Graph Interaction Model with Multi-granularity Contrastive Transfer Learning (GIMC) for zero-shot cross-lingual document-level ECI. Specifically, we introduce a heterogeneous graph interaction network to model the long-distance dependencies between events that are scattered over a document. Then, to improve cross-lingual transferability of causal knowledge learned from the source language, we propose a multi-granularity contrastive transfer learning module to align the causal representations across languages. Extensive experiments show our framework outperforms the previous state-of-the-art model by 9.4% and 8.2% of average F1 score on monolingual and multilingual scenarios respectively. Notably, in the multilingual scenario, our zero-shot framework even exceeds GPT-3.5 with few-shot learning by 24.3% in overall performance.
Keywords: zero-shot cross-lingual, document-level, event causality identification
@@
${}^{3}$ Shanghai Artificial Intelligence Laboratory, China
${}^{4}$ Ant Group, Hangzhou, China
@@ 2.2.   Heterogeneous Graph Interaction Network
After construct heterogeneous graph, we apply Graph Attention Networks v2 (GATv2) Brody et al. (2022) to model the global interactions. For each edge ($j,i$), the scoring function $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ indicates the importance of the features of the neighbor $j\in\mathcal{N}_{i}$ to the node $i$:
@@ 2.4.   Training
In our implementations, we employ the base versions of the language-specific pre-trained language models (PLMs) and the multilingual PLMs. The learning rate is initialized as 1e-3 with a linear decay. We use the AdamW algorithm Loshchilov and Hutter (2017) to optimize model parameters. The batch size is set to 1, the number of GATv2 layers is 3. The number of training epochs is 60. Each experiment is conducted on NVIDIA GeForce RTX 3090 GPUs.
@@ 3.3.   Baselines
(3) RichGCNTran Phu and Nguyen (2021) implement several interaction graphs, by learning a linear combination of the adjacency matrices of these graphs to get a final graph and uses graph convolutional network (GCN) to capture capture relevant context information.
@@ 6.   Limitations
7.   Acknowledgements
This work is supported by the National Key Research and Development Program of China (No. 2022ZD0160503), the National Natural Science Foundation of China (No. 62176257 ). This work is also supported by the Youth Innovation Promotion Association CAS.